# BubbleONet: A Physics-Informed Neural Operator for High-Frequency Bubble Dynamics

### v1

@@ Abstract
This paper introduces BubbleONet, an operator learning model designed to map pressure profiles from an input function space to corresponding bubble radius responses. BubbleONet is built upon the Physics-Informed Deep Operator Network (PI-DeepONet) framework, leveraging DeepONet’s powerful universal approximation capabilities for operator learning alongside the robust physical fidelity provided by physics-informed neural networks. To mitigate the inherent spectral bias in deep learning, BubbleONet integrates the Rowdy adaptive activation function, enabling improved representation of high-frequency features. The model is evaluated across various scenarios, including: (1) Rayleigh–Plesset equation based bubble dynamics with a single initial radius, (2) Keller–Miksis equation based bubble dynamics with a single initial radius, and (3) Keller–Miksis equation based bubble dynamics with multiple initial radii. Moreover, the performance of single-step versus two-step training techniques for BubbleONet is investigated. The results demonstrate that BubbleONet serves as a promising surrogate model for simulating bubble dynamics, offering a computationally efficient alternative to traditional numerical solvers.
@@
Physics Informed Neural Network , Bubble dynamics , Operator learning , Rayleigh–Plesset equation , Keller–Miksis equation
journal: Journal
@@ 1 Introduction
Microscopic gas bubbles are integral to a wide range of scientific and engineering applications, including medical techniques like ultrasound-mediated drug delivery [1], thermal ablation [2], and industrial operations such as cavitation-based jet cleaning [3]. On the other hand, cavitating bubbles also have harmful effects such as erosion and noise generation [4], thus making an in-depth understanding of bubbles imperative to both harnessing their benefits and mitigating their harmful effects. The dynamic behavior of these oscillating bubbles is traditionally described by models such as Rayleigh-Plesset (R-P) equation [5, 6, 7], Keller-Miksis (K-M) equation [8], Keller-Herring equation[9], and Gilmore equation [10]. These models offer essential theoretical understanding of single-bubble dynamics. However, their extension to systems involving multiple interacting bubbles introduces significant computational challenges. The resulting complexity and high computational cost often hinder their feasibility in real-time analysis and large-scale simulations.
@@
The rapid advancement of machine learning has opened new avenues for developing surrogate models that can augment or even replace traditional numerical solvers for complex physical systems; a research direction now widely recognized as scientific machine learning (SciML). Among emerging approaches, neural operator-based frameworks have garnered significant attention for their ability to learn mappings between infinite-dimensional function spaces. A leading example is the Deep Operator Network (DeepONet), introduced by Lu et al. [11], which is theoretically grounded in the universal approximation theorem for operators. The DeepONet architecture comprises of two subnetworks: a branch network that encodes the input function and a trunk network that represents the output domain. Their outputs are combined to predict the target function, enabling the model to learn nonlinear operator mappings effectively. Since its introduction, DeepONet has been widely adopted and extended across a range of scientific disciplines. For instance, Mao et al. [12] developed DeepM&Mnet to model chemical dissociation reactions, while Jin et al. [13] proposed MIOnet for multi-input operator regression, applying it to coupled diffusion–reaction systems. Peyvan et al. [14] designed a two-step training methodology to enhance DeepONet’s performance on the Riemann problem, and Goswami et al. [15] introduced a Partition-of-Unity (PoU) DeepONet framework for learning operators in chemical kinetics and turbulent flows. Additionally, Osorio et al. [16] explored the use of DeepONet for forecasting dynamics in solar-thermal energy systems. Collectively, these developments underscore DeepONet’s potential as a general-purpose tool for solving physics-informed problems through operator learning over functional spaces.
@@
Another prominent framework for solving physics-governed problems is the Physics-Informed Neural Network (PINN). The central idea behind PINNs is the incorporation of physical laws, typically expressed as partial differential equations, directly into the loss function of neural networks, thereby guiding the training process through the underlying governing equations [17]. Since their introduction by Raissi et al. [17], PINNs have been widely adopted in diverse scientific and engineering domains. By embedding physical constraints, PINNs are capable of generating highly accurate solutions while substantially reducing dependence on large labeled datasets, distinguishing them from purely data-driven models [17, 18]. To address limitations related to scalability and computational efficiency, several enhanced variants of the original PINNs framework have been developed. Notable examples include Conservative PINNs [19] and eXtended PINNs (XPINNs) [20], which leverage conservation principles and domain decomposition strategies to improve performance; see also [21, 22] for unified temporal decomposition and adaptive domain decomposition, respectively. These innovations have enabled the application of PINNs to increasingly complex and high-dimensional problems. For example, Shukla et al.  [23] employed PINNs for the identification of material properties, while Jagtap et al.  [24] addressed inverse nonlinear water wave problems. Abbasi et al. [25] developed PINN-based workflows for porous media; see also their survey on PINNs applied to shock waves [26]. Zhu et al.  [27] constructed a surrogate PINN model to simulate melt pool dynamics in metal additive manufacturing, and Mao et al.  [28] applied PINNs to the analysis of high-speed compressible fluid flows; see also Jagtap et al.  [29] for solving inverse problems in high-speed flows. Cheng et al.  [30] integrated PINNs into a representative volume element (RVE) framework to capture structure–property relationships in heterogeneous materials. Furthermore, Zhu et al.  [31] employed PINNs for surrogate modeling aimed at uncertainty quantification. PINNs have also been employed for the solution of high-dimensional partial differential equations [32, 33]. The foundational mathematical theory underlying PINNs and their extended variants (XPINNs) has been developed by several authors [34, 35, 36]. A comprehensive survey of the approximation theory for physics-informed machine learning is provided by Mhaskar et al. [37].
@@
The application of machine learning to bubble dynamics remains relatively underexplored. Lin et al. [38, 39] employed DeepONet to investigate multi-scale bubble behavior. However, their study addresses only a limited subset of the broader problem and exhibits several notable limitations. First, the pressure profiles in their simulations are generated using Gaussian random fields, which do not accurately reflect the sinusoidal nature of pressure variations encountered in various practical applications. Second, their dataset is constrained to low amplitude pressure fluctuations on the order of $10^{3}$ Pa, whereas realistic scenarios often involve amplitudes up to $10^{6}$ Pa. Additionally, a critical limitation lies in their exclusive focus on low-frequency excitations. Bubble dynamics can vary significantly under different frequency regimes, and the absence of high-frequency data impairs the model’s generalizability. These limitations emphasize the necessity for a more comprehensive machine learning framework for modeling bubble dynamics.
The objective of this study is to develop a Physics-Informed Deep Operator Network (PI-DeepONet). The proposed model, named BubbleONet, is designed to provide a robust and generalizable computational framework for the simulation of complex bubble dynamics. Wang [40] demonstrated the effectiveness of PI-DeepONet in solving a variety of partial differential equations, further reinforcing its potential for modeling intricate bubble dynamics. Goswami et al. [41] applied the PI-DeepONet to predict crack elongation in quasi-brittle materials. Lu et al. [42] employed PI-DeepONet to solve fluid dynamics problems, including diffusion and advection equations.
@@
The remainder of this paper is organized as follows. Section 2 discusses the limitations of the original DeepONet architecture in learning R-P equation-based bubble dynamics. Section 3 presents the governing equations examined in this study, the R-P equation and K-M equation. Section 4 introduces the proposed solutions to the challenges faced in previous studies, including the incorporation of adaptive activation functions and physics-based constraints. Section 5 presents the simulation results, ranging from a single-step initial radius based on R-P and K-M equations to multiple initial radius cases based on the K-M equation. Further, a comparison between the single-step training technique and the two-step training technique is also presented. To this end, Section 6 concludes the study with a summary of findings, limitations, and future directions.
2 Evaluating the Constraints of the Original DeepONet Approach to Bubble Dynamics
Lin et al. [38, 39] effectively employed this architecture to investigate bubble dynamics governed by the Rayleigh–Plesset equation. In their framework, the branch network takes the pressure variation $\Delta P$ as input, where the imposed pressure is defined as $P_{\infty}=P_{0}+\Delta P$, consistent with the assumption of small amplitude fluctuations. Here, $P_{0}$ represents the initial pressure. The trunk network receives time as its input. Both the branch and trunk networks are implemented as standard feedforward neural networks (FFNNs), and their outputs are combined through a dot product to generate the final prediction.
Parameters Range Intervals
[…]
Pressure amplitude, amp (Pa) $1,10]\times 10^{5}$ 10
Pressure frequency, $f$ (kHz) $[200,2000]$ 300
Figure 1: Sample bubble dynamics simulation obtained from APECSS for a bubble with $R_{0}=50\mu m,f=1970kHz,amp=3*10^{5}Pa,t=50\mu s$. Top: pressure profile. Bottom: Radius evolution
To evaluate the generalization capability of the model, we initially employed the original DeepONet architecture to simulate bubble dynamics. The training data were generated using the open-source code APECSS [43], which numerically solves the Rayleigh–Plesset equation. To address the limitations discussed earlier, a total of 3,000 datasets driven by sinusoidal pressure profiles were generated following the design of experiments (DoE) outlined in Table 1, with 2,400 datasets used for training and 600 for validation. An example from the dataset is shown in Figure. 1, corresponding to a bubble with an initial radius of $R_{0}=50\,\mu\text{m}$, subjected to an acoustic pressure field with a frequency of $f=1337.4\,\text{kHz}$ and an amplitude of $6\times 10^{5}\,\text{Pa}$ over a time span of $50\,\mu\text{s}$.
Figure 2: Validation results from R-P equation using the original DeepONet after 500,000 epochs. (a)Bubble dynamics driven by pressure with frequency of 630.4 kHz and amplitude of $9\times 10^{5}\,\text{Pa}$, representing low frequency case. (B)Bubble dynamics driven by pressure with frequency of 1046.8 kHz and amplitude of $3\times 10^{5}\,\text{Pa}$, representing high frequency case
Figure. 2 presents the predictions generated by the original DeepONet architecture after training. Two representative cases of pressure-driven bubble dynamics are shown, both with an initial radius of $R_{0}=50\,\mu\text{m}$. Case 1 corresponds to a pressure signal with a frequency of 630.43 kHz and an amplitude of $9\times 10^{5}\,\text{Pa}$, while Case 2 involves a relatively higher frequency of 1046.82 kHz and a lower amplitude of $3\times 10^{5}\,\text{Pa}$. To optimize model performance, we explored multiple configurations of the branch and trunk networks, varying the depth from 3 to 8 layers and the width from 128 to 512 neurons. The results displayed in Figure 2 correspond to the best-performing configuration, 8 layers with 512 neurons per layer, trained for 500,000 epochs, which yielded the lowest training and validation losses. Figure 2a illustrates the online prediction for Case 1, where the model output closely aligns with the ground truth with the low absolute error confirming the model’s ability to capture low-frequency bubble dynamics with high accuracy, where absolute error is defined as online prediction minus ground truth. In contrast, Figure 2b shows the prediction for Case 2, where the model captures the overall trend but fails to resolve the finer details of the high-frequency oscillations. This behavior is further exacerbated in cases involving ultra-high frequencies (e.g., 1900 kHz and 2000 kHz). These observations underscore the spectral bias inherent to the original DeepONet architecture. More details on the preliminary study on the original model can be found in [44].
@@ 3 Governing Equations
The physics of bubble dynamics is well characterized by mathematical models. We utilize two different equations: the Rayleigh–Plesset equation, a fundamental ordinary differential equation (ODE) that describes the idealized radial motion of a spherical bubble in an incompressible liquid [45, 6, 7], and the Keller–Miksis equation, a more comprehensive model that accounts for liquid compressibility and acoustic effects to more accurately capture bubble behavior [8].
3.1 Rayleigh-Plesset equation
@@
3.2 Keller-Miksis equation
@@
4 Methodology: A Physics-Informed Operator Learning Architecture for Simulating Bubble Dynamics
In this section, we present the architecture of BubbleONet, illustrated in Figure 3. The model comprises of two primary components: the operator layer and the operation layer. The operator layer exploits the operator approximation capabilities of the DeepONet and integrates an adaptive activation function inspired by the deep Kronecker neural network (KNN) [46]. Then, the operation layer embeds physical constraints into the learning framework to help prevent nonphysical predictions and ensure consistency with the underlying physics.
Figure 3: The architecture of BubbleONet consists of an operator layer and an operation layer. The operator layer is built on a Deep Operator Network, with a Kronecker neural network - based branch net that takes the full pressure profile as input, and a feedforward trunk net that takes time and initial radius as input. The operation layer post-processes the output to enforce physical constraints. The loss combines data and ODE components, and the optimizer updates the network parameters iteratively until convergence.
4.1 Branch net with adaptive activation function to address spectral bias
Deep learning models exhibit a well-documented phenomenon known as spectral bias, characterized by the tendency to preferentially learn simpler (low-frequency) functions over more complex (high-frequency) ones. This bias emerges not only from the training dynamics but also from the intrinsic structure and parameterization of neural networks [47, 48]. To address spectral bias, adaptive activation functions have proven effective by introducing bounded yet highly non-monotonic fluctuations into network outputs, thereby reducing saturation across layers and preserving sensitivity to high-frequency features [46]. Similar approaches are employed in the noise activation function introduced by Gulcehre [49] and the probabilistic activation function proposed by Shridhar [50]. Furthermore, adaptive activation functions enhance model performance [51] and promote smoother convergence in various deep learning frameworks, including physics-informed neural networks [52, 53]. See the comprehensive survey of activation functions by Jagtap and Karniadakis [54].
Jagtap et al.[46] introduced the Rowdy activation function, augmenting standard activation functions with sinusoidal components to effectively capture high-frequency solutions to the Helmholtz equation. This formulation was subsequently extended by incorporating a phase adjustment term, further improving the model’s capability to solve the Riemann problem[14]. Motivated by these advances, we integrate the Rowdy activation function into our BubbleONet architecture to alleviate spectral bias and enhance the network’s ability to represent high-frequency features. We begin by considering a feed-forward neural network (FFNN), expressed mathematically as follows:
$\mathbf{y}^{(l)}=\sigma^{l}\left(\mathbf{W}^{(l)}\mathbf{y}^{(l-1)}+\mathbf{b}^{(l)}\right),$ (4)
where $\mathbf{y}^{(l)}$ represents the output of layer $l$, $\mathbf{W}^{(l)}$ is the weight matrix of layer $l$, $\mathbf{y}^{(l-1)}$ is the input from the previous layer l1l-1, $\mathbf{b}^{(l)}$ is the bias vector of layer $l$, $\sigma(\cdot)$ is the activation function. Assuming a block activation function $\vec{\sigma}$ represented as:
$\vec{\sigma}(\mathbf{h})=\begin{bmatrix}\sigma_{1}(h)\\ \vdots\\ \sigma_{K}(h)\end{bmatrix}$, (5)
In our case, $\sigma_{1}$ is the ReLU activation function and $K$ is a positive integer of $5$. The remaining functions $\{\sigma_{k}\}_{k=2}^{K}$:
$\sigma(x)=n\sin\left((k-1)nx\right),\quad\forall\,2\leq k\leq K,$ (6)
Thus, using $h$ denoting the linear operation of $\mathbf{h}^{(l)}=\mathbf{W}^{(l)}\mathbf{y}^{(l-1)}+\mathbf{b}^{(l)}$, Eq 4 is written as:
$\mathbf{y}^{(l)}$ $=ReLu\left(n*a_{0}*\mathbf{h}^{(l)}\right)$ (7)
$+\sum_{i=1}^{K-1}n*a_{i}*sin\left(n*F_{i}*\mathbf{h}^{(l)}+c_{i}\right)$
where $n$ is the scaling factor set to 10 to scale the amplitude of the introduced noise, and $a_{i}$ and $c_{i}$ denote the amplitude and phase terms of the sinusoidal components. These terms are treated as learnable parameters within the neural network. $a_{i}$ and $c_{i}$ are initialized to 0, except for $a_{0}$, which is set to 0.1 to ensure that the network initially behaves similarly to a standard ReLU activation function. The Rowdy activation function is integrated into the branch network, as this component is responsible for constructing the latent representation of the input frequency-related function. Accordingly, the FFNN is replaced by a Kronecker Neural Network. Incorporating the Rowdy activation is crucial for introducing additional learnable parameters on frequencies and phases, thereby enhancing the model’s ability to capture complex bubble dynamics.

### v3

# BubbleOKAN: A Physics-Informed Interpretable Neural Operator for High-Frequency Bubble Dynamics

@@ Abstract
In this work, we employ physics-informed neural operators to map pressure profiles from an input function space to the corresponding bubble radius responses. Our approach employs a two-step DeepONet architecture. To address the intrinsic spectral bias of deep learning models, our model incorporates the Rowdy adaptive activation function, enhancing the representation of high-frequency features. Moreover, we introduce the Kolmogorov-Arnold network (KAN) based two-step DeepOKAN model, which enhances interpretability (often lacking in conventional multilayer perceptron architectures) while efficiently capturing high-frequency bubble dynamics without explicit utilization of activation functions in any form. We particularly investigate the use of spline basis functions in combination with radial basis functions (RBF) within our architecture, as they demonstrate superior performance in constructing a universal basis for approximating high-frequency bubble dynamics compared to alternative formulations. Furthermore, we emphasize on the performance bottleneck of RBF while learning the high frequency bubble dynamics and showcase the advantage of using spline basis function for the trunk network in overcoming this inherent spectral bias. The model is systematically evaluated across three representative scenarios: (1) bubble dynamics governed by the Rayleigh–Plesset equation with a single initial radius, (2) bubble dynamics governed by the Keller–Miksis equation with a single initial radius, and (3) Keller–Miksis dynamics with multiple initial radii. We also compare our results with state-of-the-art neural operators, including Fourier Neural Operators, Wavelet Neural Operators, OFormer, and Convolutional Neural Operators. Our findings demonstrate that the two-step DeepOKAN accurately captures both low- and high-frequency behaviors, and offers a promising alternative to conventional numerical solvers. The two-step DeepOKAN code is available at https://github.com/ParamIntelligence/Two-Step-DeepOKAN
@@
Bubble dynamics , Physics-Informed Neural Operator , Kolmogorov-Arnold Networks , Rayleigh–Plesset equation , Keller–Miksis equation
journal: Computer Methods in Applied Mechanics and Engineering
@@ 1 Introduction
Microscopic gas bubbles are integral to a wide range of scientific and engineering applications, including medical techniques like ultrasound-mediated drug delivery [pua2009ultrasound], thermal ablation [gnanaskandan2019modeling], and industrial operations such as cavitation-based jet cleaning [cako2022cavitation]. On the other hand, cavitating bubbles also have harmful effects such as erosion and noise generation [gnanaskandan2016large], thus making an in-depth understanding of bubbles imperative to both harnessing their benefits and mitigating their harmful effects. Understanding bubble dynamics, especially at high driving frequencies, is fundamental to a range of modern technologies. At high frequencies, bubbles experience rapid oscillations and exhibit shape instabilities that can generate intense local pressures, shear stresses, and microstreaming flows. These phenomena govern energy localization, transport, and mixing at microscales. Applications of such high-frequency bubble oscillations include ultrasound contrast imaging (where microbubbles enhance acoustic scattering), targeted drug and gene delivery (where oscillations transiently permeabilize cell membranes), and high-intensity focused ultrasound therapy (where controlled cavitation can ablate tissue). Beyond medicine, high-frequency bubble dynamics are leveraged in sonochemistry to accelerate chemical reactions, and in microfluidics for pumping, mixing, and cleaning at microscales. The dynamic behavior of these oscillating bubbles is traditionally described by models such as Rayleigh-Plesset (R-P) equation [rayleigh1917viii, franc2007rayleigh, brennen2014cavitation], Keller-Miksis (K-M) equation [keller1980bubble], Keller-Herring equation[prosperetti1986bubble], and Gilmore equation [gilmore1952growth]. These models offer essential theoretical understanding of single-bubble dynamics. However, their extension to systems involving multiple interacting bubbles introduces significant computational challenges. The resulting complexity and high computational cost often hinder their feasibility in real-time analysis and large-scale simulations.
@@
The rapid advancement of machine learning has opened new avenues for developing surrogate models that can augment or even replace traditional numerical solvers for complex physical systems; a research direction now widely recognized as scientific machine learning (SciML). Among emerging approaches, neural operator-based frameworks have garnered significant attention for their ability to learn mappings between infinite-dimensional function spaces. A leading example is the Deep Operator Network (DeepONet), introduced by Lu et al. [Lu2021], which is theoretically grounded in the universal approximation theorem for operators. The DeepONet architecture comprises of two subnetworks: a branch network that encodes the input function and a trunk network that represents the output domain. Their outputs are combined to predict the target function, enabling the model to learn nonlinear operator mappings effectively. Since its introduction, DeepONet has been widely adopted and extended across a range of scientific disciplines. For example, Mao et al. [Mao2021] developed DeepM&Mnet to model chemical dissociation reactions, while Jin et al. [Jin2022] proposed MIOnet for multi-input operator regression, applying it to coupled diffusion–reaction systems. Peyvan et al. [Peyvan2024] designed a two-step training methodology to enhance DeepONet’s performance on the Riemann problem, and Goswami et al. [goswami2024learning] introduced a Partition-of-Unity (PoU) DeepONet framework for learning operators in chemical kinetics and turbulent flows. Moreover, Osorio et al. [osorio2022forecasting] explored the use of DeepONet for forecasting dynamics in solar-thermal energy systems. Collectively, these developments highlight DeepONet’s potential as a general-purpose tool for solving physics-informed problems through operator learning over functional spaces.
@@
Another prominent SciFM methodology for solving physics-governed problems is the Physics-Informed Neural Networks (PINNs). The central idea behind PINNs is the incorporation of physical laws, typically expressed as partial differential equations, directly into the loss function of neural networks, thereby guiding the training process through the underlying governing equations [Raissi2019]. Since their introduction by Raissi et al. [Raissi2019], PINNs have been widely adopted in diverse scientific and engineering domains. By embedding physical constraints, PINNs are capable of generating highly accurate solutions while substantially reducing dependence on large labeled datasets, distinguishing them from purely data-driven models [Raissi2019, Karniadakis2021]. To address limitations related to scalability and computational efficiency, several enhanced variants of the original PINNs framework have been developed. Notable examples include Conservative PINNs [jagtap2020conservative] and eXtended PINNs (XPINNs) [jagtap2020extended], which leverage conservation principles and domain decomposition strategies to improve performance; see also [penwarden2023unified, hu2023augmented] for unified temporal decomposition and adaptive domain decomposition, respectively. These innovations have enabled the application of PINNs to increasingly complex and high-dimensional problems. For example, Shukla et al.  [shukla2021physics] employed PINNs for the identification of material properties, while Jagtap et al.  [jagtap2022deep] addressed inverse nonlinear water wave problems. Abbasi et al. [abbasi2025history] developed PINN-based workflows for porous media; see also their survey on PINNs applied to shock waves [abbasi2025challenges]. Zhu et al.  [Zhu2021] constructed a surrogate PINN model to simulate melt pool dynamics in metal additive manufacturing, and Mao et al.  [Mao2020] applied PINNs to the analysis of high-speed compressible fluid flows; see also Jagtap et al.  [jagtap2022physics] for solving inverse problems in high-speed flows. Cheng et al.  [Cheng2022] integrated PINNs into a representative volume element (RVE) framework to capture structure–property relationships in heterogeneous materials. Furthermore, Zhu et al.  [Zhu2019] employed PINNs for surrogate modeling aimed at uncertainty quantification. PINNs have also been employed for the solution of high-dimensional partial differential equations [hu2024tackling, menon2025anant]. The foundational mathematical theory underlying PINNs and their extended variants (XPINNs) has been developed by several authors [de2024error, hu2021extended, shin2020convergence]. A comprehensive survey of the approximation theory for physics-informed machine learning is provided by Mhaskar et al. [mhaskar2025approximation].
@@
The application of machine learning to bubble dynamics remains relatively underexplored. Lin et al. [Lin2021_1, Lin2021_2] employed DeepONet to investigate multi-scale bubble behavior. However, their study addresses only a limited subset of the broader problem and exhibits several notable limitations. First, the pressure profiles in their simulations are generated using Gaussian random fields, which do not accurately reflect the sinusoidal nature of pressure variations encountered in various practical applications. Second, their dataset is constrained to low amplitude pressure fluctuations on the order of $10^{3}$ Pa, whereas realistic scenarios often involve amplitudes up to $10^{6}$ Pa. Moreover, a critical limitation lies in their exclusive focus on low-frequency excitations. Bubble dynamics can vary significantly under different frequency regimes, and the absence of high-frequency data impairs the model’s generalizability. These limitations emphasize the necessity for a more comprehensive machine learning framework for modeling bubble dynamics. In this work, we overcome these limitations. The main contributions of this work are summarized as follows:
  • 1.
    We employ a physics-informed deep neural operator designed for simulating complex bubble dynamics. Specifically, we use a two-step DeepONet architecture with Rowdy activation functions [Jagtap2022]. A systematic ablation study is conducted to quantify the influence of Rowdy activations on predictive performance.
  • 2.
    Furthermore, we propose an interpretable variant based on Kolmogorov–Arnold Networks, named two-step DeepOKAN. The proposed architectures establish a robust and generalizable computational framework that integrates physical laws with deep operator learning to enhance predictive accuracy in modeling multiphase flow phenomena.
  • 3.
    We address the challenging regime of bubble dynamics characterized by high frequencies and large amplitudes, thereby extending the applicability of physics-informed operator networks to highly nonlinear fluid systems. In particular, we solve two different types of bubble dynamics equations, namely, the Rayleigh–Plesset and Keller–Miksis equations, over a large range of amplitudes ($1,10]\times 10^{5}$ Pa) and frequencies ($200,2000]\times 10^{6}$ KHz), representing a wide range of physically realistic scenarios.
  • 4.
    We provide a comprehensive benchmark for assessing the performance of physics-informed operator networks in reproducing bubble behavior under diverse flow and boundary conditions, and we compare the proposed method against state-of-the-art neural operators, including Fourier Neural Operators (FNO), Wavelet Neural Operators (WNO), OFormer, and Convolutional Neural Operators (CNO).
@@
The remainder of this paper is organized as follows. Section 2 presents the governing R-P and K-M equations for bubble dynamics. Section 3 highlights the limitations of the original DeepONet architecture in learning R-P equation-based bubble dynamics. In Section 4, we discuss various operator learning architectures, including physics-informed DeepONet, two-step DeepONet, and the proposed two-step DeepOKAN. Section 5 presents the results, ranging from single-step initial radius cases based on the R-P and K-M equations to multiple initial radius scenarios using the K-M equation with two-step DeepONet and two-step DeepOKAN. We also compare our results with other state-of-the-art methods, such as FNO, CNO, WNO, and OFormer. Furthermore, a systematic ablation study is performed with respect to the number of Rowdy activation terms. Finally, Section 6 concludes the study, summarizing the key findings, limitations, and future directions.
@@ 2 Governing Equations
The physics of bubble dynamics is well characterized by mathematical models. We utilize two different equations: the Rayleigh–Plesset equation, a fundamental ordinary differential equation (ODE) that describes the idealized radial motion of a spherical bubble in an incompressible liquid [Hong2024, franc2007rayleigh, brennen2014cavitation], and the Keller–Miksis equation, a more comprehensive model that accounts for liquid compressibility and acoustic effects to more accurately capture bubble behavior [keller1980bubble].
2.1 Rayleigh-Plesset (R-P) equation
@@
2.2 Keller-Miksis (K-M) equation
@@
3 Limitations of the Vanilla DeepONet Approach to Bubble Dynamics
Lin et al. [Lin2021_1, Lin2021_2] employed the Vanilla DeepONet architecture to investigate bubble dynamics governed by the Rayleigh–Plesset equation. In their framework, the branch network takes the pressure variation $\Delta P$ as input, where the imposed pressure is defined as $P_{\infty}=P_{0}+\Delta P$, consistent with the assumption of small amplitude fluctuations. Here, $P_{0}$ represents the initial pressure. The trunk network receives time as its input. Both the branch and trunk networks are implemented as standard feedforward neural networks (FFNNs), and their outputs are combined through a dot product to generate the final prediction.
Parameters Range #\# Samples
[…]
Pressure amplitude, amp (Pa) $1,10]\times 10^{5}$ 10
Pressure frequency, $f$ (KHz) $[200,2000]$ 300
Figure 1: A bubble dynamics simulation based on R-P equation obtained from APECSS for a bubble with $R_{0}=50\mu m,f=1970\text{KHz},~\text{amp}=3\times 10^{5}Pa,t=50\mu s$. Top: pressure profile. Bottom: Radius evolution
To evaluate the generalization capability of the model, we initially employed this vanilla DeepONet architecture to simulate bubble dynamics. The training data were generated using the open-source code APECSS [Denner2023], which numerically solves the Rayleigh–Plesset equation. A total of 3,000 datasets driven by sinusoidal pressure profiles were generated following the design of experiments (DoE) outlined in Table 1, with 2,400 datasets used for training and 600 for validation. An example from the dataset is shown in Figure 1, corresponding to a bubble with an initial radius of $R_{0}=50\,\mu\text{m}$, subjected to an acoustic pressure field with a frequency of $f=1970\,\text{KHz}$ and an amplitude of $3\times 10^{5}\,\text{Pa}$ over a time span of $50\,\mu\text{s}$.
Figure 2: Validation results from R-P equation using the vanilla DeepONet after 500,000 epochs. (a) Bubble dynamics driven by pressure with frequency of 676 KHz and amplitude of $3\times 10^{5}\,\text{Pa}$, representing low frequency case. (b) Bubble dynamics driven by pressure with frequency of 1654 KHz and amplitude of $10\times 10^{5}\,\text{Pa}$, representing high frequency case.
Figure 2 presents the predictions generated by the vanilla DeepONet architecture after training. Two representative cases of pressure-driven bubble dynamics are shown, both with an initial radius of $R_{0}=50\,\mu\text{m}$. Case 1 corresponds to a pressure signal with a frequency of 630.43 KHz and an amplitude of $9\times 10^{5}\,\text{Pa}$, while Case 2 involves a relatively higher frequency of 1654 KHz and a lower amplitude of $10\times 10^{5}\,\text{Pa}$. To optimize model performance, we explored multiple configurations of the branch and trunk networks, varying the depth from 3 to 8 layers and the width from 128 to 512 neurons. The results displayed in Figure 2 correspond to the best-performing configuration, 8 layers with 512 neurons per layer, trained for 500,000 epochs, which yielded the lowest training and validation losses. Figure 2(a) illustrates the online prediction for Case 1, where the model output closely aligns with the ground truth, with a low absolute error confirming the model’s ability to capture low-frequency bubble dynamics with high accuracy. In contrast, Figure 2(b) shows the prediction for Case 2, where the model captures the overall trend but fails to resolve the finer details of the high-frequency oscillations. This behavior is further exacerbated in cases involving ultra-high frequencies (e.g., 1900 KHz and 2000 KHz). These observations highlight the spectral bias (a propensity to learn low-frequency components more readily than high-frequency ones) inherent to the original DeepONet architecture. More details on the preliminary study on the original model can be found in [zhang2025bubble].
4 Operator Learning Architectures for High-Frequency Bubble Dynamics
Among the various neural operator architectures, DeepONets have shown remarkable flexibility and efficiency across a range of physical systems. In this section, we first discuss the physics-informed DeepONet architecture [wang2021learning], which integrates governing physical laws directly into the training process to enhance data efficiency and enforce physical consistency. We then discuss the two-step DeepONet architecture [lee2024training], which is more accurate and efficient than vanilla DeepONet. To this end, we propose a more interpretable version based on Kolmogorov-Arnold Networks (KAN), named as, two-step DeepOKAN architecture.
4.1 Single-Step DeepONet
Figures 3 depict the schematic of the physics-informed DeepONet architecture, wherein physical constraints are seamlessly embedded alongside the vanilla DeepONet framework to enforce governing laws.
Figure 3: Architecture of DeepONet: The operator layer employs a Kronecker-based branch net (input: full pressure profile) and a trunk net (input: time and initial radius). The operation layer enforces physical constraints on outputs, while training minimizes a combined data-ODE loss via iterative optimization.
Deep learning models are known to exhibit spectral bias arising from both network parameterization and training dynamics [Cao2021, Rahaman2019]. Adaptive activation functions have been proposed to mitigate this bias by introducing bounded, non-monotonic variations that reduce saturation and enhance sensitivity to high-frequency features [Jagtap2022]. Related concepts include the noise activation function [Gulcehre2016] and probabilistic activation function [Shridhar2019], both of which improve model expressivity. Such adaptive activations also promote smoother convergence and improved accuracy across diverse architectures, see, [Jagtap2020, jagtap2020locally], for more details; a comprehensive overview is provided by Jagtap and Karniadakis [jagtap2023important]. In particular, Jagtap et al. [Jagtap2022] proposed the Rowdy activation function, which augments conventional activation functions with sinusoidal terms to capture high-frequency behaviors in problems such as the Helmholtz equation and the Riemann problem featuring discontinuities [Peyvan2024]. Building on these advances, the present work integrates the Rowdy activation within the branch network of the DeepONet architecture. This design introduces learnable frequency and phase parameters, thereby alleviating spectral bias and enhancing the model’s capacity to represent complex bubble dynamics.